OpenStudy (anonymous):

X-2Y= 2 X*Y=12 Me ajudem nessa.

OpenStudy (perl):

(1) x*y= 12 (2) x -2y = 2 from (1) y = 12/x substitute into (2) X - 2(12/X) = 2 x - 24/x = 2 x^2 - 24 = 2x x^2 -2x - 24 = 0 x = 6,-4

OpenStudy (anonymous):

'-' nossa obg

OpenStudy (perl):

what language is that?

OpenStudy (anonymous):

portuguese , thank you